OpenStudy (anonymous):

FInd the derivative of the function by using the definition of the derivative… please help f(x)=sqrtx + 1

OpenStudy (anonymous):

it is all inside the radical? \[f(x)=\sqrt{x+1}\]?

OpenStudy (anonymous):

noo only x is inside..

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[f(x)=\sqrt{x}+1\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[f'(x)=\lim_{h\to 0}\frac{f(x+h)-f(x)}{h}\] so we get \[\frac{\sqrt{x+h}+1-(\sqrt{x}+1)}{h}\] and now some algebra

OpenStudy (anonymous):

evidently the ones up top add up to zero, so we can ignore them and compute \[\frac{\sqrt{x+h}-\sqrt{x}}{h}\] gimmick is to multiply top and bottom by the conjugate of the numerator, namely \(\sqrt{x+h}+\sqrt{x}\) leave the denominator in factored form you get \[\frac{(\sqrt{x+h}-\sqrt{x})(\sqrt{x+h}+\sqrt{x})}{h(\sqrt{x+h}+\sqrt{x})}\] \[=\frac{x+h-x}{h(\sqrt{x+h}+\sqrt{x})}\]

OpenStudy (anonymous):

all that is left in the numerator is \(h\) which cancels with the \(h\) in the denominator leaving \[\frac{1}{\sqrt{x+h}+\sqrt{x}}\] now take the limit by replacing \(h\) by zero and you are done

OpenStudy (anonymous):

I got lost on the third part… :/ @satellite73

OpenStudy (anonymous):

this one \[\frac{(\sqrt{x+h}-\sqrt{x})(\sqrt{x+h}+\sqrt{x})}{h(\sqrt{x+h}+\sqrt{x})}\] ??

OpenStudy (anonymous):

yes that one,,… but I think I kinda know what you did… thats the product rule right???

OpenStudy (anonymous):

ok it is this, you want to get rid of the radical, so you know \((a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

which means \[(\sqrt{h+h}-\sqrt{x})(\sqrt{x+h}+\sqrt{x})=\sqrt{x+h}^2-\sqrt{x}^2\] \[=x+h-x=h\]

OpenStudy (anonymous):

yess...

OpenStudy (anonymous):

but isn't that a rule….

OpenStudy (anonymous):

and your real job is to figure out how to get rid of the \(h\) in the denominator

OpenStudy (anonymous):

so now that you have an \(h\) up top and a factor of \(h\) in the bottom, with \[\frac{h}{h(\sqrt{x+h}+\sqrt{x})}\] you can cancel the \(h\) giving \[\frac{1}{\sqrt{x+h}+\sqrt{x}}\] now you can replace \(h\) by 0 without getting a zero in the denominator

OpenStudy (anonymous):

ok and to do that you replace it with zero..?

OpenStudy (anonymous):

yes, now you can take the limit as \(h\to 0\) by simply erasing the \(h\)

OpenStudy (anonymous):

ahhh ok ok… makes sense.

OpenStudy (anonymous):

write \[\lim_{h\to 0}\frac{1}{\sqrt{x+h}+\sqrt{x}}=\frac{1}{\sqrt{x}+\sqrt{x}}\] \[=\frac{1}{2\sqrt{x}}\]

OpenStudy (anonymous):

and dont forget this one the derivative of \[\sqrt{x}\] is always \[\frac{1}{2\sqrt{x}}\] so while your classmates are trying to figure it out on a test or quiz using the stupid exponential notation and the power rule, you write the answer instantly. it never changes, you never have to compute it again

OpenStudy (anonymous):

omg thankyou so much. we have a quiz on this tomorrow also. thank you . really . Im glad you actually went through and took the time to explain this to me . thank you again.

OpenStudy (anonymous):

@satellite73